# Machines are more productive than humans until they aren’t, and vice versa

### v1

@@ Abstract
With the growth of artificial skills, organizations may increasingly confront with the problem of optimizing skill policy decisions guided by economic principles. This paper addresses the underlying complexity of this challenge by developing an in-silico framework based on Monte Carlo simulations grounded in empirical realism to analyze the economic impact of human and machine skills, individually or jointly deployed, in the execution of tasks presenting varying levels of complexity. Our results provide quantitative support for the established notions that automation tends to be the most economically-effective strategy for tasks characterized by low-to-medium generalization difficulty, while automation struggles to match the economic utility of human skills in more complex scenarios. Critically, our simulations highlight that combining human and machine skills can be the most effective strategy when a high level of generalization is required, but only if genuine augmentation is achieved. In contrast, when failing to realize this synergy, the human-machine policy is severely penalized by the inherent costs of its dual skill structure, causing it to destroy value and becoming the worst choice from an economic perspective. The takeaway for decision-makers is unambiguous: simply allocating human and machine skills to a task is insufficient, and a human-machine skill policy is neither a silver-bullet solution nor a low-risk compromise. Rather, it is a critical opportunity to boost competitiveness that demands a strong organizational commitment to enabling augmentation. Also, our findings show that improving the cost-effectiveness of machine skills over time, while useful, does not replace the fundamental need to focus on achieving augmentation. By synthesizing task-based economic theory with the concept of generalization difficulty, our work offers a scalable and low-cost methodology for organizations to forecast the economic implications of their choices. This provides a foundation for testing skill policies ex ante and for scaling experiments to study more general policy implications as candidates for future empirical validation.
@@ 4.3 Margin of contribution and value
It is worth noticing that the generality of the $\delta_{c}$ parameters comes at the price of a simplification, because no specific assumption about the factors influencing the organization efficiency (or productivity) are made. For example, as illustrated in [Farboodi and Veldkamp, 2021], learning production techniques from data is one path that organizations can walk to increase their efficiency, and with data being a by-product of the organization’s activity, the more productive the firm is, the more data it can generate and store, thus entering the so-called data feedback loop. Naturally, modeling this effect goes far beyond setting the slope of a linear growth. The design we propose through the $\delta_{c}$ parameters represents a first step, while a non-linear function of cumulative data or experience may be a promising direction for future developments.
@@ 5 Simulation model
By leveraging the properties of the model illustrated in the previous Section 4, we deploy the method of Monte Carlo simulation to produce synthetic data. The scope of this section is two-folded: in a first part, we illustrate the logics governing the simulation, and we formalize the description of the simulation output serving as input for the subsequent analyses; in a second part, we discuss the challenges underlying the use of simulation as a tool for observing emerging phenomena, detailing the reasons why we think that this approach can b viable and valuable for our research.
@@ 5.2 Rationale for the Monte Carlo method
A fundamental critique of computational simulation is that its results risk to be a mere consequence of the hypotheses and assumptions encoded in the model (the tautology critique), unless appropriate reasons justify its use. We believe that acknowledging this risk is a necessary step, together with a detailed rationale, to support the choice of simulation in the use case of interest.
@@ 8 Conclusions
The simulation’s initial conditions are set to establish a clear baseline, with Beta distributions modeling the performance of a specific and representative scenario. Initially, the machine skill is superior on low-to-medium difficulty tasks, reflecting its high initial efficiency. Over time, the human skill evolves, partially closing the performance gap on low and medium-difficulty tasks, while simultaneously demonstrating a widening advantage on high-difficulty tasks. The margin of contribution for automation is set to be empirically higher than for human labor to reflect lower variable costs and higher economies of scale. Human-machine interaction, initially floating in a range compatible with the sum of the costs of both human and machine skills, improves its margin of contribution over epochs up to a factor of 2x. No constraint is set for the combinations of the minimum performance triggering an error shock and its unitary cost. This setup, combined with non-linear performance-to-output relationships, contribute to characterize an experimental scenario that is grounded in empirical pragmatism, suitable to representing a common real-world case.
@@
Moreover, our results quantitatively support the notions that, while automation is most beneficial for low-to-medium difficulty tasks, it is insufficient for high-difficulty scenarios where the cost of machine errors can erode all economic utility generated by high unitary margins of contribution. Then, the second key takeaway for decision-makers pertains the importance of knowing as better as possible the process of interest, and the characteristics of the environment in which the task unfolds. Being able to estimate the characteristics of the probability distribution of task difficulty can be an advantage in the making of skill policy decisions that can affect the cost structure of the organization.
@@ A.4 From data to skill policies
In a more general setting, at any arbitrary time $t$ characterized by the state of technological advancement $T$, given the task of interest $z_{i}$ and its description $D_{i,t}$, the decision-maker confronts with the following problem:
@@
where $E\big[U({Sol}_{i,t,c,j})\big]$ is the expected utility achieved by deploying the skill ${Sol}_{i,t,c,j}$, where $j$ denotes a specific skill. For example, in case of a heuristic ($c=H$), $j$ may represent different procedures, or different operators; in contrast, for statistical models ($c=M$), $j$ represents different and alternative classes of models (a neural network, or random forest, or a linear regression), as well as alternative implementations of the same model.
@@
In Figure 29 we illustrate a schematic representation of the process, generalizing to the case of perturbations of the training/evaluation set (for example, by deploying cross-validation). As we see, the distributions $P(\theta)_{c[,e]}$ serving as input for the simulation process do not come from thin air, but they can be estimated through a rigorous process that may involve multiple candidate skill program instances. Additionally, in Figure 30 we describe the additional complexity of estimating $P(\theta)_{c,d[,e]}$ when it is possible to decompose $E_{i,t,c[,q]}$ by task difficulty $d$. In this representation, we consider producing a model that is trained on data representing all the task difficulties, but if the decomposition of situational examples is done on the full data set $D_{i,t,c}$, then it may be possible to create dedicated models for each level of task difficulty (this choice depends on the actual possibility to sort task by difficulty at inference time, which may not be feasible for a variety of reasons).
@@
This general treatment illustrates a procedural path from data to skill policy. However, while formally straightforward, a real-world practical example may contextualize even better the effort.
@@ A.4.1 A practical example
The organization of interest is a real-estate agency, whose business is based on commissions (3% of the transaction price). A core decision-making problem for the agency is to determine the optimal price for properties, such that the total commissions earned in a year is maximized. A price lower than the optimal one makes a sale faster, but reduces the satisfaction of the seller (the primary agency’s customer) and reduces the commission, which is computed as a percentage of the transaction price. In contrast, a price higher than the optimal one makes the sale slower and more difficult, increasing the total resources allocated by the agency to each specific sale opportunity and potentially reducing the number of transaction in the period of interest. As a result, the definition of an optimal price is a crucial factor for the agency to achieve economic health and wealth. Additionally, as the turnover of sales representatives is increasingly high, creating a proprietary machine skill that becomes a distinctive agency asset is seen as strategic from the agency’s shareholders. The mean cost of error is estimated as 10% of the MAPE (the true price for a property is computed by adjusting the actual transaction price by the lead time of sale, such that the true price tends to be higher than the true potential transaction price for quick sales, and tends to be lower in case of slow sales (the average transaction price in the last 3 years is 200k and the average number of sales per year is 150). In this context, the market of technological solutions is growing, and the agency has been offered two alternative solutions (A and B). Now, the decision-making problem faced by the agency is two-folded: first, deciding whether to decline both the models A and B, or introducing one of them; then, in case of the introduction of one or more model(s), defining the most economically-effective skill policy: automation of the prediction task (machine-exclusive skill policy), or human-machine interaction?
@@
Heuristics used by experienced sales agents, with knowledge of the market (both offer and demand). In this case, we consider two categories of agents: senior (more than 5 years in the local business, or more than 100 sales) with MAPE = 15(%), or junior with MAPE = 20.
@@
Model A is an on-premise solution based on a linear regression, with MAPE = 10 and a development fee of 50k (negligible operating costs); model B is a cloud-based solution implementing an ensemble of regressors (linear, random forest, and neural network) with MAPE = 2, no development costs, licensed for 25k/year.
@@
The agency plans to expanding its activity to new areas, with emerging new high-profile constructions but lower connections with the city center and fewer services. It is opinion of the agency director that predicting the optimal price for these new properties is more complex and difficult for sales representatives, and especially for the senior ones (they may be biased). In contrast, junior reps may be more effective in their predictions, with lower bias with respect to senior ones, but with the tendency to show higher variability of judgment.

### v2

@@ Abstract
With the growth of artificial skills, organizations may increasingly confront with the problem of optimizing skill policy decisions guided by economic principles. This paper addresses the underlying complexity of this challenge by developing an in-silico framework based on Monte Carlo simulations grounded in empirical realism to analyze the economic impact of human and machine skills, individually or jointly deployed, in the execution of tasks presenting varying levels of complexity. Our results provide quantitative support for the established notions that automation tends to be the most economically-effective strategy for tasks characterized by low-to-medium generalization difficulty, while automation may struggle to match the economic utility of human skills in more complex scenarios. Critically, our simulations highlight that combining human and machine skills can be the most effective strategy when a high level of generalization is required, but only if genuine augmentation is achieved. In contrast, when failing to realize this synergy, the human-machine policy is severely penalized by the inherent costs of its dual skill structure, causing it to destroy value and becoming the worst choice from an economic perspective. The takeaway for decision-makers is unambiguous: in contexts requiring high generalization capabilities, simply allocating human and machine skills to a task is insufficient, and a human-machine skill policy is neither a silver-bullet solution nor a low-risk compromise. Rather, it is a critical opportunity to boost competitiveness that demands a strong organizational commitment to enabling augmentation. Also, our findings show that improving the cost-effectiveness of machine skills over time, while useful, does not replace the fundamental need to focus on achieving augmentation. By synthesizing task-based economic theory with the concept of generalization difficulty, our work offers a scalable and low-cost methodology for organizations to forecast the economic implications of their choices. This provides a foundation for testing skill policies ex ante and for scaling experiments to study more general policy implications as candidates for future empirical validation.
@@ 4.3 Margin of contribution and value
It is worth noticing that the generality of the $\delta_{c}$ parameters comes at the price of a simplification, because no specific assumption about the factors influencing the organization efficiency (or productivity) is made. For example, as illustrated in [Farboodi and Veldkamp, 2021], learning production techniques from data is one path that organizations can walk to increase their efficiency, and with data being a by-product of the organization’s activity, the more productive the firm is, the more data it can generate and store, thus entering the so-called data feedback loop. Naturally, modeling this effect goes far beyond setting the slope of a linear growth. The design we propose through the $\delta_{c}$ parameters represents a first step, while a non-linear function of cumulative data or experience may be a promising direction for future developments.
@@ 5 Simulation model
By leveraging the properties of the model illustrated in the previous Section 4, we deploy the method of Monte Carlo simulation to produce synthetic data. The scope of this section is two-folded: in a first part, we illustrate the logics governing the simulation, and we formalize the description of the simulation output serving as input for the subsequent analyses; in a second part, we discuss the challenges underlying the use of simulation as a tool for observing emerging phenomena, detailing the reasons why we think that this approach can be viable and valuable for our research.
@@ 5.2 Rationale for the Monte Carlo method
A fundamental critique of computational simulation is that its results risk to be a mere consequence of the hypotheses and assumptions encoded in the model (the tautology critique), unless reasons justify its use. We believe that acknowledging this risk is a necessary step, together with a detailed rationale, to support the choice of simulation in the use case of interest.
@@ 8 Conclusions
The simulation’s setup conditions are set to establish a clear baseline, with Beta-distributed probabilities modeling a scenario where the state of technological advancement for machine skills is not yet able to provide consistent and widespread performance when confronting with tasks requiring high levels of generalization difficulty. More specifically, initially, the machine skill is superior on low-to-medium difficulty tasks, reflecting its high initial efficiency. Over time, the human skill evolves, partially closing the performance gap on low and medium-difficulty tasks, while simultaneously demonstrating a widening advantage on high-difficulty tasks. The margin of contribution for automation is set to be empirically higher than for human labor to reflect lower variable costs and higher economies of scale. Human-machine interaction, initially floating in a range compatible with the sum of the costs of both human and machine skills, improves its margin of contribution over epochs up to a factor of 2x. No constraint is set for the combinations of the minimum performance triggering an error shock and its unitary cost. This setup, combined with non-linear performance-to-output relationships, contribute to characterize an experimental scenario that is grounded in empirical pragmatism, suitable to representing a common real-world case.
@@
Moreover, our results quantitatively support the notions that, while automation is most beneficial for low-to-medium difficulty tasks, it may be insufficient for high-difficulty scenarios where the cost of machine errors can erode all economic utility generated by high unitary margins of contribution. This conclusion is contingent upon the initial assumptions pertaining to the probability distributions of performance for human and machine skills. In our framework, machine skills are not omniscient, and high generalization difficulty represents their fundamental limitation. Should advancements successfully address this Achilles heel, the economic rationale against automating high-difficulty tasks may dissolve, opening new frontiers for value creation.
Also, a second key takeaway for decision-makers pertains the importance of knowing as better as possible the process of interest, and the characteristics of the environment in which the task unfolds. Being able to estimate the characteristics of the probability distribution of task difficulty can be an advantage in the making of skill policy decisions that can affect the cost structure of the organization.
@@ A.4 From data to skill policies
In a more general setting, at any arbitrary time $t$ characterized by a specific state of technological advancement, given the task of interest $z_{i}$ and its description $D_{i,t}$, the decision-maker confronts with the following problem:
@@
where $E\big[U({Sol}_{i,t,c,j})\big]$ is the expected utility achieved by deploying the skill ${Sol}_{i,t,c,j}$, where $j$ denotes a specific implementation (or instance) of a skill [^2] The notational characterization of a generic technological solution as $Sol$ is derived from [Chollet, 2019], in which $Sol_{T}^{\theta}$ denotes the shortest solution for the task $T$ achieving performance $\theta$ at evaluation time.. For example, in case of a heuristic ($c=H$), $j$ may represent different procedures, or different operators; in contrast, for statistical models ($c=M$), $j$ represents different and alternative classes of models (a neural network, or random forest, or a linear regression), as well as alternative implementations of the same model.
@@
In Figure 29 we illustrate a schematic representation of the process, generalizing to the case of perturbations of the training/evaluation set (for example, by deploying cross-validation). As we see, the distributions $P(\theta)_{c[,e]}$ serving as input for the simulation process do not come from thin air, but they can be estimated through a rigorous process that may involve multiple candidate skill program instances. In Figure 30 we describe the additional complexity of estimating $P(\theta)_{c,d[,e]}$ when it is possible to decompose $E_{i,t,c[,q]}$ by task difficulty $d$. In this representation, we consider producing a model that is trained on data representing all the task difficulties, but if the decomposition of situational examples is done on the full data set $D_{i,t,c}$, then it may be possible to create dedicated models for each level of task difficulty (this choice depends on the actual possibility to sort task by difficulty at inference time, which may not be feasible for a variety of reasons).
@@
This general treatment illustrates a procedural path from data to skill policy. However, while formally straightforward, an example inspired by a practical use case may contextualize even better the effort.
@@ A.4.1 A practical example
To explore the logics and utility of task difficulty and human-machine augmentation as factors playing a role in the analysis of a skill policy decision problem, we walk through a practical example. It should be noted that the data presented in the following example is hypothetical and illustrative, not empirical, and has been constructed solely for clarifying purpose.
@@
The fictitious organization of interest for our example is a real-estate agency, whose business is based on commissions (3% of the transaction price). A core decision-making problem for the agency is to determine the optimal price for properties, such that the total commissions earned in a year is maximized. A price lower than the optimal one makes a sale faster, but reduces the satisfaction of the seller (the primary agency’s customer) and reduces the commission, which is computed as a percentage of the transaction price. In contrast, a price higher than the optimal one makes the sale slower and more difficult, increasing the total resources allocated by the agency to each specific sale opportunity and potentially reducing the number of transaction in the period of interest. As a result, the definition of an optimal price is a crucial factor for the agency to achieve economic health and wealth. Additionally, as the turnover of sales representatives is increasingly high, creating a proprietary machine skill that becomes a distinctive agency asset is seen as strategic from the agency’s shareholders. The mean cost of error is estimated as 10% of the MAPE (the true price for a property is computed by adjusting the actual transaction price by the lead time of sale, such that the true price tends to be higher than the true potential transaction price for quick sales, and tends to be lower in case of slow sales (the average transaction price in the last 3 years is 200k and the average number of sales per year is 150). In this context, the market of technological solutions is growing, and the agency has been offered two alternative solutions (A and B). Now, the decision-making problem faced by the agency is two-folded: first, deciding whether to decline both the models A and B, or introducing one of them; then, in case of the introduction of one or more model(s), defining the most economically-effective skill policy: automation of the prediction task (machine-exclusive skill policy), or human-machine interaction?
@@
Heuristics used by experienced sales agents, with knowledge of the market (both offer and demand). In this case, we consider two categories of agents: Senior (more than 5 years in the local business, or more than 100 sales), or Junior.
@@
Model A is an on-premise solution based on a linear regression, with a development fee of 200k and negligible operating costs. Model B is a cloud-based solution implementing an ensemble of regressors (linear, random forest, and neural network) with 50k development costs, licensed for 100k/year.
@@
The agency plans to expanding its activity to new areas, with emerging new high-profile constructions but lower connections with the city center and fewer services. It is opinion of the agency director that predicting the prices in these new area may be more complex, because the target customer is different from the typical one know by the agency (then, higher levels of generalization difficulty are required in this context). Simplifying, we denote the prediction of prices for properties located in the typical area served by the agency as a low-difficulty task, while a prediction related to properties located in the new areas is considered a high-difficulty task. It is also opinion of the agency director that Senior sales representatives may be more effective in the former, while Junior reps may be more effective in the latter. To validate this conjecture, the agency director commissions an empirical analysis aimed at measuring the Mean Absolute Percentage Error (MAPE) of all the candidate predictive skills. The MAPE measured through the analysis is the following: 15% for Senior sales representative, 25% for the Junior ones, and 15% for both Model A and Model B.